OpenStudy (anonymous):

Multiply and simplify:

OpenStudy (anonymous):

The problem is: \[(3+\sqrt{6})(5+\sqrt{6})\]

OpenStudy (lilg132):

\[21 + 8\sqrt{6}\]

OpenStudy (anonymous):

The answer choices are: A. \[21+8\sqrt{6}\] B. \[15+2\sqrt{6}\] C. \[18+5\sqrt{6}\]

OpenStudy (anonymous):

:)